Please help me with this problem asap thankspreview answers Remaining time: 32:47 (min:sec) Problem 13. (20 points) Suppose a bag contains many marbles, 49% of which are
Answer #1

Solution

Given that ,

a) p = 0.49

1 - p = 1 - 0.49 = 0.51

n = 5

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2 ) = (5 C 2) * (0.49)2 * (0.51)3

= 0.3184

Probability = 0.318

b) expected value = n * p = 5 * 0.49 = 2.45